A piece of equipment will function only all components A, B and C are working. The probability of failing during one year is 15%, that of B failing in 5% & that of C failing in 10%. What is the probability that the equipment will fail before the end of the year?

Grade:12

1 Answers

chandan munny
13 years ago

dear subham gupta

taking the probability of a failing as 15%

let

the probablity of a failing =p(a)=0.15

the probablity of b failing =p(b)=0.05

the probablity of c failing =p(c)=0.10

the probability that a wll fail in one year is (0.15*0.95*0.90)+(0.85*0.05*0.90)+(0.85*0.95*0.10)